# Robust Predictive Motion Planning by Learning Obstacle Uncertainty

### v1

@@ Abstract
Safe motion planning for robotic systems in dynamic environments is nontrivial in the presence of uncertain obstacles, where estimation of obstacle uncertainties is crucial in predicting future motions of dynamic obstacles. The worst-case characterization gives a conservative uncertainty prediction and may result in infeasible motion planning for the ego robotic system. In this paper, an efficient, robust, and safe motion-planing algorithm is developed by learning the obstacle uncertainties online. More specifically, the unknown yet intended control set of obstacles is efficiently computed by solving a linear programming problem. The learned control set is used to compute forward reachable sets of obstacles that are less conservative than the worst-case prediction. Based on the forward prediction, a robust model predictive controller is designed to compute a safe reference trajectory for the ego robotic system that remains outside the reachable sets of obstacles over the prediction horizon. The method is applied to a car-like mobile robot in both simulations and hardware experiments to demonstrate its effectiveness.
@@ II-A Uncertainty Prediction
The robust approach for predicting uncertainties of a dynamic system is by formulating a sequence of sets over the horizon to cover the worst-case uncertainty realizations. This is typically achieved by forward reachability analysis [8]. For example, [2] characterized the forward reachable sets (FRSs) of the ego robotic system considering the worst-case disturbances for safety guarantees in robust motion planning. In [9], the FRSs were predicted for both an autonomous ego vehicle and a surrounding vehicle based on the maximal disturbance set for safety verification. Focusing on uncertainty predictions of other traffic participants, the later work in [10] and [11] combined the worst-case FRSs with the road networks and the interactions with the autonomous ego vehicle, respectively, in practical driving scenarios. An alternative yet more conservative approach to the forward reachability analysis is computing the minimal robust positive invariant (MRPI) set for the ego system under the maximal disturbances, and relevant works can be found in [12], [13], and [14].
@@ III-A Notations
$\mathbb{R}^{n}$ is the $n$-dimensional real number vector space, $\mathbb{R}_{+}^{n}$ is the $n$-dimensional non-negative vector space, $\mathbb{N}^{n}$ is the $n$-dimensional natural number vector space, and $\mathbb{N}_{+}^{n}$ is the $n$-dimensional positive integer vector space. $[0,1]^{n}$ means an $n$-dimensional vector with each element between $[0,1]$. $I^{n}$ indicates an $n\times n$ identity matrix. Matrices of appropriate dimension with all elements equal to 1 and 0 are denoted by $\bm{1}$ and $\bm{0}$, respectively. An interval of integers is denoted by $\mathbb{I}_{a}^{b}=\{a,a+1,\cdots,b\}$. For two sets $\mathcal{A}$ and $\mathcal{B}$, $\mathcal{A}\oplus\mathcal{B}=\left\{x+y\ |\ x\in\mathcal{A},y\in\mathcal{B}\right\}$. For a set $\Gamma\subset\mathbb{R}^{n_{a}+n_{b}}$, the set projection is defined as $\mathcal{A}={\rm Proj}_{a}(\Gamma)=\left\{a\in\mathbb{R}^{n_{a}}:\exists b\in\mathbb{R}^{n_{b}},(a,b)\in\Gamma\right\}$. For the MPC iteration, the current time step is indicated by $t$. A prediction of a variable $x$ at time step $t+i$ over the prediction horizon is represented as $x_{i|t}$, where $i=1,\ldots,N$, and $N$ is the prediction horizon. The sampling interval is represented as $T$. The cardinality of a set $\mathcal{A}$ is represented by $|\mathcal{A}|$.
@@ III-B Modeling of Ego System and Surrounding Obstacles
where $A^{s}_{t}$ and $B^{s}_{t}$ are the system matrices at time step $t$, $x^{s}_{t}\in\mathbb{R}^{n^{s}_{x}}$ is the state and $u^{s}_{t}\in\mathbb{R}^{n^{s}_{u}}$ is the input at time step $t$. The superscript $s$ indicates the $s$-th obstacle and the index set of all involved obstacles is denoted by a finite set $\mathcal{S}$.
@@
The LTV model (2) is used for several reasons. First, a linear model can easily be derived to approximate the motion characteristics of the obstacle. Second, linear models for the obstacles facilitate efficient computation of forward reachable sets when predicting the motion of obstacles.
@@ IV-A Learning Intended Control Set $\tilde{\mathbb{U}}_{t}^{s}$
At the current time step $t$, the ego system can measure or infer the control input $u^{s}_{t-1}$ of model (2) of the obstacle $s$, $\forall s\in\mathcal{S}$.
The estimation of the control input $u_{t-1}^{s}$ depends on observed information of the obstacle $s$ and the specific form of model (2). This will be specified in Remark VI.2 in the case studies. Since the intended control set is a hidden variable, an approach is to infer the hidden information using past observations [32]. To this end, based on Assumption IV.1, we define the information set $\mathcal{I}^{s}_{t}$ available to the ego system at time step $t$, which consists of the collected inputs of the $s$-th obstacle until time step $t$. The information set is updated as
@@
$\mathbb{U}^{s}(v,\rho,\bm{\theta})$ $=f(H^{s},\bm{\theta})\oplus(1-\rho)v,$ (6a)
@@
$\mathop{\rm minimize}\limits_{v,\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (7)
${\rm subject\ to}\quad$ $H^{s}u^{s}-(1-\rho)H^{s}v\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
@@
$\mathop{\rm minimize}\limits_{y,\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (8)
${\rm subject\ to}\quad$ $H^{s}u^{s}-H^{s}y\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
@@
$\mathop{\rm minimize}\limits_{\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (10)
${\rm subject\ to}\quad$ $H^{s}u^{s}\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
@@
$\mathop{\rm minimize}\limits_{y,\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (11)
${\rm subject\ to}\quad$ $H^{s}u^{s}-H^{s}y\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
@@
Note that in (11), the constraint $H^{s}y\leq\bm{0}$ implies that $y=\bm{0}$. This confirms that (10) and (11) are equivalent with $\rho=1$. Hence, we conclude that (7) and (8) are equivalent, and the optimal quantified set $\hat{\mathbb{U}}_{t}^{s}$ is given in the form of (9). ∎
@@
The formulation (8) is a standard randomized program [33], and it is shown in [33, Theorem 1] that the probability of intersection between $\hat{\mathbb{U}}_{t}^{s}$ and $\tilde{{\mathbb{U}}}_{t}^{s}$ is positively related to $|\mathcal{I}_{t}^{s}|$ if $\hat{\mathbb{U}}_{t}^{s}$ is invariant and the control action satisfies independent and identical distribution. While these assumptions are generally stringent in practice, the results highlight that the robustness of $\hat{\mathbb{U}}_{t}^{s}$ is enhanced with a larger information set $\mathcal{I}_{t}^{s}$.
@@ IV-B2 Online Recursion
$\mathop{\rm minimize}\limits_{v,\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (12)
${\rm subject\ to}\quad$ $u_{t-1}^{s}\in{\mathbb{U}}^{s}(v,\rho,\bm{\theta}),$
@@
$\mathop{\rm minimize}\limits_{y,\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (13)
${\rm subject\ to}\quad$ $H^{s}u_{t-1}^{s}-H^{s}y\leq\bm{\theta},$
@@
where $y^{\star}_{\rm pre}$ and $\bm{\theta}^{\star}_{\rm pre}$ are the optimal solution of (13) at time step $t-1$, i.e., $\hat{\mathbb{U}}^{s}_{t-1}=f(H^{s},\bm{\theta}_{\rm pre}^{\star})\oplus y_{\rm pre}^{\star}$. The formulation (13) is computationally efficient as the number of constraints does not change with $|\mathcal{I}_{t}^{s}|$, and it does not lose any information when new samples are observed and added in the set $\mathcal{I}_{t}^{s}$.
@@ IV-C Forward Reachability Analysis Based on $\hat{\mathbb{U}}_{t}^{s}$
Given $\hat{\mathbb{U}}_{t}^{s}$ obtained by (13), and the system described as in (2), forward reachability can be used to predict the occupancy of obstacles. Denote by $\hat{\mathcal{O}}_{i|t}^{s}$ the occupancy of the center of geometry of obstacle $s$ predicted $i$ steps ahead of the current time step $t$. The set $\hat{\mathcal{O}}_{i|t}^{s}$ can be computed as[4]
@@
Figure 3: The learned intended control sets by formulation (8), formulation (13), the moving-horizon (MH) method, and the rigid tube (RT) method (for the MH method $|\mathcal{I}_{t}^{s}|=30$ ). (a) The number of samples is $50$. (b). The number of samples is $500$.
@@
Methods Computation time
Number of samples = $50$ Number of samples = $500$
Formulation (13) $2.5\ {\rm ms}$ $2.6\ {\rm ms}$
Formulation (8) $61.34\ {\rm ms}$ $695.8\ {\rm ms}$
Moving-horizon method $30.94\ {\rm ms}$ $28.9\ {\rm ms}$
Rigid tube method $1.6\ {\rm ms}$ $1.2\ {\rm ms}$

### v2

@@ Abstract
Safe motion planning for robotic systems in dynamic environments is nontrivial in the presence of uncertain obstacles, where estimation of obstacle uncertainties is crucial in predicting future motions of dynamic obstacles. The worst-case characterization gives a conservative uncertainty prediction and may result in infeasible motion planning for the ego robotic system. In this paper, an efficient, robust, and safe motion-planning algorithm is developed by learning the obstacle uncertainties online. More specifically, the unknown yet intended control set of obstacles is efficiently computed by solving a linear programming problem. The learned control set is used to compute forward reachable sets of obstacles that are less conservative than the worst-case prediction. Based on the forward prediction, a robust model predictive controller is designed to compute a safe reference trajectory for the ego robotic system that remains outside the reachable sets of obstacles over the prediction horizon. The method is applied to a car-like mobile robot in both simulations and hardware experiments to demonstrate its effectiveness.
@@ II-A Uncertainty Prediction
The robust approach for predicting uncertainties of a dynamic system is by formulating a sequence of sets over the horizon to cover the worst-case uncertainty realizations. This is typically achieved by forward reachability analysis [8], [9]. For example, [2] characterized the forward reachable sets (FRSs) of the ego robotic system considering the worst-case disturbances for safety guarantees in robust motion planning. In [10], the FRSs were predicted for both an autonomous ego vehicle and a surrounding vehicle based on the maximal disturbance set for safety verification. Focusing on uncertainty predictions of other traffic participants, the later work in [11] and [12] combined the worst-case FRSs with the road networks and the interactions with the autonomous ego vehicle, respectively, in practical driving scenarios. An alternative yet more conservative approach to the forward reachability analysis is computing the minimal robust positive invariant (MRPI) set for the ego system under the maximal disturbances, and relevant works can be found in [13], [14], and [15].
@@ III-A Notations
$\mathbb{R}^{n}$ is the $n$-dimensional real number vector space, $\mathbb{R}_{+}^{n}$ is the $n$-dimensional non-negative vector space, $\mathbb{N}^{n}$ is the $n$-dimensional natural number vector space, and $\mathbb{N}_{+}^{n}$ is the $n$-dimensional positive integer vector space. $[0,1]^{n}$ means an $n$-dimensional vector with each element between $[0,1]$. $I^{n}$ indicates an $n\times n$ identity matrix. Matrices of appropriate dimension with all elements equal to 1 and 0 are denoted by $\bm{1}$ and $\bm{0}$, respectively. An interval of integers is denoted by $\mathbb{I}_{a}^{b}=\{a,a+1,\cdots,b\}$. For two sets $\mathcal{A}$ and $\mathcal{B}$, $\mathcal{A}\oplus\mathcal{B}=\left\{x+y\ |\ x\in\mathcal{A},y\in\mathcal{B}\right\}$. For a set $\Gamma\subset\mathbb{R}^{n_{a}+n_{b}}$, the set projection is defined as $\mathcal{A}={\rm Proj}_{a}(\Gamma)=\left\{a\in\mathbb{R}^{n_{a}}:\exists b\in\mathbb{R}^{n_{b}},(a,b)\in\Gamma\right\}$. For the MPC iteration, the current time step is indicated by $t$. A prediction of a variable $x$ at time step $t+i$ over the prediction horizon is represented as $x_{i|t}$, where $i=1,\ldots,N$, and $N$ is the prediction horizon. The sampling interval is represented as $T$. The cardinality of a set $\mathcal{A}$ is represented by $|\mathcal{A}|$. The weighted inner product of a vector $x$ is denoted by $||x||_{A}^{2}=x^{\top}Ax$.
@@ III-B Modeling of Ego System and Surrounding Obstacles
where $A^{s}_{t}$ and $B^{s}_{t}$ are the system matrices at time step $t$, which are determined according to specific kinematic or dynamic equations of the obstacle. The vector $x^{s}_{t}\in\mathbb{R}^{n^{s}_{x}}$ is the state and $u^{s}_{t}\in\mathbb{R}^{n^{s}_{u}}$ is the input at time step $t$. The superscript $s$ indicates the $s$-th obstacle and the index set of all involved obstacles is denoted by a finite set $\mathcal{S}$.
@@
The LTV model (2) is used for two main reasons. (1) Since the true dynamics of the obstacle are unknown, a linear model can easily be derived to approximate the motion characteristics of the obstacle. It benefits the computation of forward reachable sets when predicting the motion of obstacles. (2) The linear model also facilitates the estimation of the control inputs of obstacles for the proposed method, while this can be challenging or computationally expensive for nonlinear models.
@@ IV-A Learning Intended Control Set $\tilde{\mathbb{U}}_{t}^{s}$
The system (2) does not have redundant actuators, and at the current time step $t$, the state $x_{t}^{s}$ can be measured or estimated by the ego system.
This implies that the matrix $B_{t-1}^{s}$ of model (2) has full column rank, such that at the current time step $t$, the ego system can find unique solution when solving for $u_{t-1}^{s}$, i.e., the control input of (2) at the previous time step $t-1$, from the expression $B_{t-1}^{s}u_{t-1}^{s}=x_{t}^{s}-A_{t-1}^{s}x_{t-1}^{s}$. For example, the solution $u_{t-1}^{s}$ in (4) is $u_{t-1}^{s}=[(v_{x,t}^{s}-v_{x,t-1}^{s})/T\ (v_{y,t}^{s}-v_{y,t-1}^{s})/T]^{\top}$. Since the intended control set is a hidden variable, an approach is to infer the hidden information using past observations [33]. To this end, based on Assumption IV.1, we define the information set $\mathcal{I}^{s}_{t}$ available to the ego system at time step $t$, which consists of the collected inputs of the $s$-th obstacle until time step $t$. The information set is updated as
@@
$\mathbb{U}^{s}(v,\rho,\bm{\theta})$ $=f(H^{s},\bm{\theta})\oplus\{(1-\rho)v\},$ (6a)
@@
$\mathop{\rm minimize}\limits_{v,\rho,\bm{\theta}}$ $\bm{1}^{\top}\bm{\theta}+\rho$ (7)
${\rm subject\ to}$ $H^{s}u^{s}-(1-\rho)H^{s}v\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
@@
$\mathop{\rm minimize}\limits_{y,\rho,\bm{\theta}}$ $\bm{1}^{\top}\bm{\theta}+\rho$ (8)
${\rm subject\ to}$ $H^{s}u^{s}-H^{s}y\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
@@
$\mathop{\rm minimize}\limits_{\rho,\bm{\theta}}$ $\bm{1}^{\top}\bm{\theta}+\rho$ (10)
${\rm subject\ to}$ $H^{s}u^{s}\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
@@
$\mathop{\rm minimize}\limits_{y,\rho,\bm{\theta}}$ $\bm{1}^{\top}\bm{\theta}+\rho$ (11)
${\rm subject\ to}$ $H^{s}u^{s}-H^{s}y\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
@@
The following shows that the constraint $H^{s}y\leq\bm{0}$ in (11) is satisfied if and only if $y=\bm{0}$.
Since the origin is an interior of the set $\mathbb{U}^{s}$, suppose $H^{s}y\leq\bm{0}$ and $y\neq\bm{0}$ hold. This means $\forall\varepsilon\in\mathbb{R}_{+}$ such that $H^{s}(\varepsilon y)\leq\bm{0}\leq\bm{1}$, this further implies that $\varepsilon y\in\mathbb{U}^{s}$, where $\varepsilon y$ is the terminal of a line segment starting from the origin. However, since the set $\mathbb{U}^{s}$ is bounded, increasing the value of $\varepsilon$ will eventually cause $\varepsilon y$ to exceed the limit of $\mathbb{U}^{s}$. Therefore, there does not exist any $y\neq{\bm{0}}$ satisfying $H^{s}y\leq{\bm{0}}$, and the only feasible solution of the variable $y$ in (11) is $y=\bm{0}$.
This confirms that (10) and (11) are equivalent with $\rho=1$. Hence, we conclude that (7) and (8) are equivalent, and the optimal quantified set $\hat{\mathbb{U}}_{t}^{s}$ is given in the form of (9). ∎
@@
Formulation (8) is a standard randomized program [34]. It is shown in [22, Theorem 3] that if $\tilde{\mathbb{U}}_{t}^{s}$ is time-invariant, the sequence of control actions are independent and identically distributed, and $|\mathcal{I}_{t}^{s}|$ is large enough, then with high confidence, the set $\hat{\mathbb{U}}_{t}^{s}$ is a good approximation of the true intended set $\tilde{{\mathbb{U}}}_{t}^{s}$. While these assumptions are stringent in practice, the results highlight that the robustness of $\hat{\mathbb{U}}_{t}^{s}$ is enhanced with a larger information set $\mathcal{I}_{t}^{s}$.
@@ IV-B2 Online Recursion
$\mathop{\rm minimize}\limits_{v,\rho,\bm{\theta}}$ $\bm{1}^{\top}\bm{\theta}+\rho$ (12)
${\rm subject\ to}$ $u_{t-1}^{s}\in{\mathbb{U}}^{s}(v,\rho,\bm{\theta}),$
@@
$\mathop{\rm minimize}\limits_{y,\rho,\bm{\theta}}$ $\bm{1}^{\top}\bm{\theta}+\rho$ (13)
${\rm subject\ to}$ $H^{s}u_{t-1}^{s}-H^{s}y\leq\bm{\theta},$
@@
where $y^{\star}_{\rm pre}$ and $\bm{\theta}^{\star}_{\rm pre}$ are the optimal solution of (13) at time step $t-1$, i.e., $\hat{\mathbb{U}}^{s}_{t-1}=f(H^{s},\bm{\theta}_{\rm pre}^{\star})\oplus y_{\rm pre}^{\star}$. The formulation (13) is computationally efficient as the number of constraints does not change with $|\mathcal{I}_{t}^{s}|$, and it includes all samples in the set $\mathcal{I}_{t}^{s}$. The online recursion (13) is initialized with $\hat{\mathbb{U}}_{0}^{s}$, which is obtained by solving (8) with an initial information set $\mathcal{I}_{0}^{s}$. The set can be formulated as a small, yet nonempty, set with artificial samples.
The characteristics of the moving-horizon approach and online recursion approach will be demonstrated by a set of numerical examples in Section IV-D.
@@ IV-C Forward Reachability Analysis Based on $\hat{\mathbb{U}}_{t}^{s}$
Given $\hat{\mathbb{U}}_{t}^{s}$ obtained in Section IV-B, and the system described as in (2), forward reachability can be used to predict the obstacle occupancy. Denote by $\hat{\mathcal{O}}_{i|t}^{s}$ the occupancy of the center of geometry of obstacle $s$ predicted $i$ steps ahead of the current time step $t$. The set $\hat{\mathcal{O}}_{i|t}^{s}$ can be computed as[4]
@@
Figure 3: The learned intended control sets by formulation (8), formulation (13), the moving-horizon (MH) method, and the rigid tube (RT) method (for the MH method $L=30$ ). (a) The number of samples is $50$, $\mathbb{U}^{s}$ is a square. (b) The number of samples is $500$, $\mathbb{U}^{s}$ is a square. (c) The number of samples is $50$, $\mathbb{U}^{s}$ is a regular hexagon. (d) The number of samples is $500$, $\mathbb{U}^{s}$ is a regular hexagon. (e) The set volume with colors according to Fig. 3(a). The dashed lines indicate the results with $\mathbb{U}^{s}$ as a regular hexagon, and solid lines indicate the results with $\mathbb{U}^{s}$ as a square.
@@
Methods Formulation (8) Formulation (13) Moving Horizon Rigid Tube Convex Hull
Learn Set Reach. Pred. Learn Set Reach. Pred. Learn Set Reach. Pred. Learn Set Reach. Pred. Learn Set Reach. Pred. Vertices
$50$ Samples, $\mathbb{U}^{s}$ as a square 8.70 ${\rm ms}$ 36.3 ${\rm ms}$ 1.80 ${\rm ms}$ 34.8 ${\rm ms}$ 5.09 ${\rm ms}$ 32.4 ${\rm ms}$ 1.45 ${\rm ms}$ 32.0 ${\rm ms}$ 0.766 ${\rm ms}$ 1138 ${\rm ms}$ 18
$50$ Samples, $\mathbb{U}^{s}$ as a hexagon 12.7 ${\rm ms}$ 498 ${\rm ms}$ 2.38 ${\rm ms}$ 544 ${\rm ms}$ 6.84 ${\rm ms}$ 500 ${\rm ms}$ 1.47 ${\rm ms}$ 531 ${\rm ms}$ 0.867 ${\rm ms}$ 2477 ${\rm ms}$ 22
$500$ Samples, $\mathbb{U}^{s}$ as a square 40.1 ${\rm ms}$ 38.4 ${\rm ms}$ 1.61 ${\rm ms}$ 35.8 ${\rm ms}$ 4.92 ${\rm ms}$ 34.8 ${\rm ms}$ 1.34 ${\rm ms}$ 34.2 ${\rm ms}$ 0.770 ${\rm ms}$ 23476 ${\rm ms}$ 42
$500$ Samples, $\mathbb{U}^{s}$ as a hexagon 101 ${\rm ms}$ 514 ${\rm ms}$ 2.19 ${\rm ms}$ 560 ${\rm ms}$ 6.97 ${\rm ms}$ 568 ${\rm ms}$ 1.35 ${\rm ms}$ 555 ${\rm ms}$ 0.737 ${\rm ms}$ 12043 ${\rm ms}$ 34